# CQ-DINO: Mitigating Gradient Dilution via Category Queries for Vast Vocabulary Object Detection

### v1

@@ 4 Experiments
4.1 Datasets
We conduct experiments on two detection benchmarks:
(1) V3Det [37]: A vast vocabulary detection dataset containing 13,204 categories with 183k training and 30k validation images. This dataset represents our primary benchmark for evaluating vast vocabulary detection capabilities. (2) COCO val2017 [19]: A standard benchmark dataset with 80 common object categories comprising 118k training and 5k validation images. We include this dataset to verify the effectiveness of our method in limited vocabulary scenarios.
4.2 Implementation details
All experiments are conducted on 8 NVIDIA A100-40G GPUs with a total batch size of 16. We follow baseline methods’ original configurations for fair comparison. We implement three Swin Transformer [23] variants as backbones: Swin-B (ImageNet-1k [7] pre-trained), Swin-B-22k (ImageNet-22k pre-trained), and Swin-L (ImageNet-22k pre-trained). Our category queries are initialized using text embeddings from CLIP-ViT-L [26]. To handle dataset-specific vocabulary variations, we employ 100 category queries for V3Det and 30 queries for COCO, aligned with their respective category sizes. We derive the hierarchical tree structure from V3Det’s category taxonomy for tree construction. For the implicit relation learning, we employ a standard 8-head self-attention module.
The training objective combines multiple loss terms with the following weights: classification loss (Asymmetric loss [29], weight=1.0), contrastive alignment (Focal loss [30], weight=1.0), bounding box regression (L1 loss, weight=5.0), and GIoU loss [28] (weight=2.0), as in GroundingDINO [22]. We employ Hungarian matching following GroundingDINO [22] with identical matching costs for the object-to-query assignment. To stabilize training, we use a two-stage training approach. In the first stage, we pre-train the category queries, image encoder, and image-guided query selection module for 10 epochs. This creates a high initial recall of target categories before fine-tuning the complete detection pipeline in the second stage.
Method Epochs Backbone $AP$ $AP_{50}$ $AP_{75}$ ATSS  [47] 24 Swin-B 7.6 8.9 8.0 FCOS [35] 24 Swin-B 21.0 24.8 22.3 Faster R-CNN [27] 24 Swin-B 37.6 46.0 41.1 CenterNet2 [52] 24 Swin-B 39.8 46.1 42.4 Cascade R-CNN [2] 24 Swin-B 42.5 49.1 44.9 Deformable DETR [54] 50 Swin-B 42.5 48.3 44.7 DINO [46] 24 Swin-B 42.0 46.8 43.9 Prova [4] 24 Swin-B 44.5 49.9 46.6 CQ-DINO (Ours) 24 Swin-B 46.3 51.5 48.4 DINO [46] 24 Swin-B-22k 43.4 48.4 45.4 Prova [4] 24 Swin-B-22k 50.3 56.1 52.6 CQ-DINO (Ours) 24 Swin-B-22k 52.3 57.7 54.6 DINO [46] 24 Swin-L 48.5 54.3 50.7 Prova [4] 24 Swin-L 50.9 57.2 53.2 CQ-DINO (Ours) 24 Swin-L 53.0 58.4 55.4
@@ 4.2 Implementation details
Table 2: Comparison with state-of-the-art DETR variants on COCO val2017 (best results reported from original papers).
@@ 4.3 Experimental results
Performance on standard detection benchmark. Tab. 2 compares our method with state-of-the-art approaches on COCO val2017. Despite being primarily designed for vast vocabulary scenarios, CQ-DINO achieves competitive performance, reaching 58.5% AP, comparable with recent DETR-based methods. To ensure a fair comparison, we report the best results in their original papers. Results demonstrate that our method achieves gradient rebalancing and adaptive hard mining as in limited vocabulary scenarios.
ECC IGQS $AP$ $AR^{C}_{100}$ FPS
Out Of Memory
TC Out Of Memory
51.1 80.9 6.7
SA 51.3 ($\uparrow 0.2$) 75.5 ($\downarrow 5.4$) 6.5 ($\downarrow 0.2$)
TC 52.3 ($\uparrow 1.2$) 83.3 ($\uparrow 2.4$) 6.6 ($\uparrow 0.1$)
Table 3: Ablation study on the effectiveness of encoding category correlations (ECC) and image-guided query selection (IGQS) components in CQ-DINO on V3Det dataset using Swin-Base-22k backbone. SA denotes Self-Attention module and TC refers to hierarchical tree construction. Metrics include AP, category-level recall ($AR^{C}_{100}$), and inference speed (FPS). The absence of IGQS results in CUDA memory overflow.
@@ 4.4 Ablation studies
We conduct comprehensive ablation studies to analyze the effectiveness of each component in the proposed CQ-DINO. All experiments are performed on the V3Det dataset using a Swin-B-22k backbone unless otherwise specified.
Effect of each component in CQ-DINO. Tab. 3 presents component contributions using $AP$ and $AR_{100}^{C}$ (category-level average recall with 100 selected queries).
Image-guided query selection resolves memory constraints. Without this component, the baseline fails to run on A100-40G GPUs even with batch size 1 due to high category processing demands.
The category query framework shows superior flexibility over conventional detection heads in modeling inter-category relationships. Tree construction delivers gains of 1.2% AP and 2.4% $AR_{100}^{C}$ through explicit hierarchical relationship modeling. self-attention’s implicit approach achieves marginal +0.2% AP but reduces $AR_{100}^{C}$. This stems from difficulties learning complex relationships across 13k+ categories. Tree construction introduces zero parameters with only 0.1 FPS overhead, while self-attention adds 2.36M parameters and 0.2 FPS overhead.
While Tab. 3 shows that explicit tree construction outperforms self-attention on V3Det’s vast category space, we conduct further experiments on COCO (Table 6). Results reveal that self-attention contributes meaningful 0.2% AP and 0.9% $AR_{30}^{C}$ improvements on datasets with fewer categories. Both experiments validate the effectiveness of encoding category correlations, with the optimal approach depending on the scale of the category space.
Query Count $AP$ $AR^{C}$ FPS
50 51.8 78.4 6.9
100 52.3 83.3 6.6
200 52.2 87.5 5.9
Method $AP$ $AP_{50}$ $AP_{75}$ $AR^{C}_{100}$
Fixed weight (0.5) 51.9 57.4 54.3 82.3
Ours ($\alpha_{v}$) 52.3 57.7 55.4 83.3
@@
Method $AP$ $AR_{30}^{C}$ Params (M)
CQ-DINO w/o SA 58.3 98.2 244.3
CQ-DINO w/ SA 58.5 ($\uparrow 0.2$) 99.1 ($\uparrow 0.9$) 246.7 ($+2.4$)
Impact of category query count. Tab. 4 examines how varying the number of category queries (50/100/200) affects performance. As expected, increasing the number of queries improves $AR_{100}^{C}$ at the cost of lower inference speed. However, simply maximizing query count is not optimal. With too many queries, the gradient imbalance between positive and negative examples becomes problematic. Our experiments show that 100 queries achieve optimal balance between detection and computational efficiency for V3Det. For COCO dataset experiments, 30 queries are sufficient, achieving 99.1% $AR_{30}^{C}$.
Effectiveness of adaptive weighting in tree construction. Tab. 5 shows the importance of our adaptive weighting strategy compared to a fixed weight of 0.5. This approach adjusts weights based on the varying number of child categories for each parent node in the hierarchy. Results show our adaptive approach outperforms fixed weighting.
Method Params/Cat. Memory/Cat. Max Cats.
(K) (kB) (k)
DINO [46] 2.1 8.9 100
CQ-DINO 0.8 2.7 130
@@
4.5 Discussion
Scalability of CQ-DINO. Tab. 7 analyzes the scaling efficiency of CQ-DINO compared to DINO for vast vocabulary detection. Our method requires only 0.8K parameters per category, representing a 62% reduction from DINO’s 2.1K parameters. For runtime memory consumption, CQ-DINO uses 2.7KB CUDA memory per category. To evaluate practical scalability limits, we test the maximum category support on an A100-40G GPU using Swin-B-22k backbone with a single $800\times 1333$ resolution image. CQ-DINO supports detection of up to 130k categories, surpassing DINO’s 100k limit. Experiments validate that CQ-DINO enables applications with extremely larger vocabularies.
Method Backbone $AP$ $AP_{50}$ $AP_{75}$ GenerateU [18] Swin-L&T5-B 0.4 0.5 0.4 ChatRex [15] Swin-L&ChatRex-7B 1.3 1.5 1.4 GenerateU* [18] Swin-L&T5-B 17.2 22.4 17.7 DINO [46] Swin-L 48.5 54.4 50.7 CQ-DINO (Ours) Swin-L 53.0 58.4 55.4
Table 8: Performance of generated-based methods on V3Det dataset. We report zero-shot performance using their most powerful models. The line marked with * * represents GenerateU finetuned on the V3Det dataset.
Limitations of generation-based methods. We evaluate generation-based methods on the V3Det dataset in Tab. 8. Following the evaluation protocol from GenerateU [18], we compute semantic similarity between generated category embeddings and V3Det category embeddings, selecting the highest similarity match as the final prediction. Our experiments reveal poor performance: GenerateU [18] achieves only 0.4% AP, while the more recent ChatRex [15] performs just 1.3% AP. This highlights a fundamental limitation: generation-based methods struggle to control the granularity of generated categories, creating semantic misalignments with the specific requirements of detection tasks. Furthermore, even when we finetune GenerateU on V3Det data, the resulting performance (17.2% AP) still exhibits a substantial gap compared to classification-head methods.
$\gamma$ $\alpha$ 0.25 0.35 0.50 0.75
2 43.4 45.1 47.4
3 43.7 43.9 45.1
5
@@ 4.5 Discussion
Focal loss parameter analysis. As analyzed in Sec. 3.1, vast vocabulary detection suffers from gradient dilution challenges. Focal loss (FL) [30] mitigates this via adaptive weighting with hyperparameters $\alpha$ and $\gamma$. Theoretically, $\alpha$ balances positive/negative sample contributions, where increasing $\alpha$ enhances the model’s ability to handle more categories. Factor $\gamma$ focuses learning on hard negatives, where increasing $\gamma$ improves the mining of hard examples.
Tab. 9 evaluates FL configurations on DINO with a Swin-B-22k backbone. The baseline [37] ($\alpha=0.25$, $\gamma=2$) achieves 43.4% AP. We observe training instability for $\gamma\geq 5$ or $\alpha\geq 0.75$. Optimal performance (47.2% AP, +4.0% over baseline) occurs at $\gamma=2$ and $\alpha=0.5$, confirming that proper parameter tuning can substantially benefit vast vocabulary detection.
Interestingly, when examining parameter transferability across architectures in Tab. 10, we find that this optimal setting does not generalize well. The same parameters ($\alpha=0.5$, $\gamma=2$) lead to performance degradation of 3.3% AP for Swin-B backbone compared to default configuration ($\alpha=0.25$, $\gamma=2$). However, they increase the performance by 4.0% AP and 1.6% AP for Swin-B-22k and Swin-L backbones, respectively. These findings show that while Focal loss addresses gradient dilution issues, its optimal configuration requires careful parameter tuning.
4.6 Limitation
While CQ-DINO achieves improvement in vast vocabulary object detection, several limitations need future exploration. First, our detection performance depends on the recall capability of our first-stage category query selection. The detector cannot recover them later if relevant categories are not included in the top-K selected queries. Fortunately, our approach achieves 83.3% AR, ensuring this component does not constitute a performance bottleneck in practice. Second, the two-stage training paradigm introduces additional complexity to the training pipeline. This separate optimization process could potentially limit the performance.
5 Conclusion
In this work, we analyze the challenges in vast vocabulary detection: positive gradient dilution and hard negative gradient dilution. To address these limitations, we propose CQ-DINO with two core innovations: (1) learnable category queries that encode category correlations, and (2) image-guided query selection that reduces negative space while performing adaptive hard negative mining. CQ-DINO demonstrates superior performance on the V3Det benchmark and exhibits strong scalability to larger vocabularies. In the future, we will explore the flexibility of category queries for open-vocabulary detection and incremental learning scenarios.

### v3

@@ 4 Experiments
4.1 Datasets and Implementation Details
We conduct experiments on two detection benchmarks: (1) V3Det [35]: a vast vocabulary detection dataset containing 13,204 categories, with 183k training and 30k validation images. This is our primary benchmark for evaluating vast vocabulary detection. (2) COCO val2017 [19]: a standard benchmark dataset with 80 object categories, comprising 118k training and 5k validation images. We include this dataset to verify the effectiveness of our method in limited vocabulary scenarios.
@@ 4.1 Datasets and Implementation Details
Method Epochs Backbone $AP$ $AP_{50}$ $AP_{75}$
ATSS  [42] 24 Swin-B 7.6 8.9 8.0
FCOS [33] 24 Swin-B 21.0 24.8 22.3
Faster R-CNN [27] 24 Swin-B 37.6 46.0 41.1
CenterNet2 [46] 24 Swin-B 39.8 46.1 42.4
Cascade R-CNN [2] 24 Swin-B 42.5 49.1 44.9
Deformable DETR [47] 50 Swin-B 42.5 48.3 44.7
DINO [41] 24 Swin-B 42.0 46.8 43.9
Prova [4] 24 Swin-B 44.5 49.9 46.6
CQ-DINO (Ours) 24 Swin-B 46.3 51.5 48.4
DINO [41] 24 Swin-B-22k 43.4 48.4 45.4
Prova [4] 24 Swin-B-22k 50.3 56.1 52.6
CQ-DINO (Ours) 24 Swin-B-22k 52.3 57.7 54.6
DINO [41] 24 Swin-L 48.5 54.3 50.7
Prova [4] 24 Swin-L 50.9 57.2 53.2
CQ-DINO (Ours) 24 Swin-L 53.0 58.4 55.4
Table 2: Comparison between the proposed CQ-DINO and state-of-the-art DETR variants on COCO val2017, reporting the best results as provided by their respective original papers.
@@
Table 3: Performance of Open-world methods on V3Det dataset. We report zero-shot performance using their strongest models. * * indicates finetuned results.
Method Backbone $AP$ $AP_{50}$ $AP_{75}$ GenerateU [18] Swin-L&T5-B 0.4 0.5 0.4 ChatRex [14] Swin-L&LLM-7B 1.3 1.5 1.4 GenerateU* [18] Swin-L&T5-B 21.8 27.2 22.1 DINO [41] Swin-L 48.5 54.4 50.7 CQ-DINO (Ours) Swin-L 53.0 58.4 55.4
Experiments are conducted on 8 A100-40G GPUs with a total batch size of 16, unless otherwise specified. Baseline configurations are used for fair comparison. Three Swin Transformer [23] variants serve as backbones: Swin-B (ImageNet-1k [7] pre-trained), Swin-B-22k (ImageNet-22k pre-trained), and Swin-L (ImageNet-22k pre-trained). Category queries are initialized using CLIP-ViT-L [26] text embeddings. To match dataset vocabularies, we employ 100 category queries for V3Det and 30 for COCO, aligned with their category sizes. The hierarchical tree structure is derived from V3Det’s category taxonomy. For implicit relation learning, we use an 8-head self-attention module.
The training objective combines multiple loss terms with the following weights: classification loss (Asymmetric Loss [29], weight=1.0), contrastive alignment (Focal Loss [30], weight=1.0), bounding box regression (L1 Loss, weight=5.0), and GIoU Loss [28] (weight=2.0), as in GroundingDINO [22]. Hungarian matching is used, following GroundingDINO, with identical matching costs for object-to-query assignment. To stabilize training, we use a two-stage approach: first, pre-training category queries, image encoder, and image-guided query selection for 10 epochs to establish high initial target category recall; then, fine-tuning the full detection pipeline in the second stage.
@@ 4.2 Experimental Results
Performance on Standard Detection Benchmark. Tab. 3 compares our method with state-of-the-art approaches on COCO val2017. Despite being primarily designed for vast vocabulary scenarios, CQ-DINO achieves competitive performance, reaching 58.5% AP, which is comparable to recent DETR-based methods. We report the best results from their original papers to ensure a fair comparison. The competitive performance of CQ-DINO is mainly due to the proposed gradient rebalancing and adaptive hard mining strategies.
Table 4: Ablation study on the effectiveness of encoding category correlations and image-guided query selection components in CQ-DINO on V3Det dataset with Swin-Base-22k backbone. “–” denotes unavailable $AR^{C}$ metrics due to absence of query selection.
Encoding Category Correlations Image-guided Query Selection $AP$ $AR^{C}$ FPS
47.3 0.7
Hierarchical Tree construction 49.4 ($\uparrow 2.1$) 0.6 ($\downarrow 0.1$)
51.1 80.9 10.8
Self-Attention module 51.3 ($\uparrow 0.2$) 75.5 ($\downarrow 5.4$) 10.4 ($\downarrow 0.4$)
Hierarchical Tree Construction 52.3 ($\uparrow 1.2$) 83.3 ($\uparrow 2.4$) 10.6 ($\downarrow 0.2$)
@@ 4.3 Ablation Studies
We conduct ablation studies to evaluate the effectiveness of each component in CQ-DINO. Unless otherwise stated, all experiments are performed on the V3Det dataset using a Swin-B-22k backbone.
Effect of Each Component in CQ-DINO. Table 4 presents the contribution of each component in terms of average precision ($AP$) and category-level average recall with selected queries ($AR^{C}$). To establish a baseline without image-guided query selection, we conduct experiments on 8 H800-80G GPUs due to memory constraints. Notably, incorporating image-guided query selection increases FPS from 0.7 to 10.8, highlighting its effectiveness in alleviating memory bottlenecks and substantially enhancing inference efficiency. Furthermore, this component addresses the issue of gradient dilution and leads to a substantial AP improvement, from 47.3% to 51.1% in detection performance.
Explicit hierarchical modeling via tree construction leads to an improvement of 1.2% AP and 2.4% $AR^{C}$. In contrast, employing self-attention for implicit relationship modeling achieves a marginal increase of 0.2% AP but reduces $AR^{C}$. This is due to difficulties in learning complex relationships across 13k+ categories. Notably, the tree construction method introduces zero additional parameters with only a 0.2 FPS overhead, while the self-attention approach adds 2.36M extra parameters and a 0.4 FPS reduction in inference speed.
While Tab. 4 shows that explicit tree construction outperforms self-attention on V3Det’s vast category space, we conduct further experiments on COCO (Table 6). The results reveal that self-attention contributes meaningful 0.2% AP and 0.9% $AR^{C}$ improvements on datasets with fewer categories. Both experiments validate the effectiveness of encoding category correlations, with the optimal approach depending on the scale of the category space.
Method $AP$ $AR^{C}$ Params (M) CQ-DINO w/o SA 58.3 98.2 244.3 CQ-DINO w/ SA 58.5 ($\uparrow 0.2$) 99.1 ($\uparrow 0.9$) 246.7 ($+2.4$)
@@
Method $AP$ $AP_{50}$ $AP_{75}$ $AR^{C}$ Fixed weight (0.5) 51.9 57.4 54.3 82.3 Ours ($\alpha_{v}$) 52.3 57.7 55.4 83.3
Effectiveness of Adaptive Weighting in Tree Construction. Tab. 6 shows the importance of our adaptive weighting strategy compared to a fixed weight of 0.5. This approach adjusts weights based on the varying number of child categories for each parent node in the hierarchy. Results show that our adaptive approach outperforms fixed weighting.
4.4 Discussion
Scalability of CQ-DINO. Tab. 8 compares the scaling efficiency of CQ-DINO and DINO for vast vocabulary detection. CQ-DINO requires only 0.8K parameters per category, representing a 62% reduction from DINO’s 2.1K parameters. For runtime memory consumption, CQ-DINO uses 2.7KB CUDA memory per category. To evaluate practical scalability limits, we test the maximum category support on an A100-40G GPU using the Swin-B-22k backbone with a single $800\times 1333$ resolution input image. CQ-DINO supports detection of up to 130k categories, surpassing DINO’s 100k limit. These experiments validate that CQ-DINO enables applications with extremely large vocabularies.
@@ 4.4 Discussion
Method Params/Cat. Memory/Cat. Max Cats. (K) (kB) (k) DINO [41] 2.1 8.9 100 CQ-DINO 0.8 2.7 130
@@
$\gamma$ $\alpha$ 0.25 0.35 0.50 0.75 2 43.4 45.1 47.4 3 43.7 43.9 45.1 5
Limitations of Generation-based Methods. We evaluate generation-based methods on the V3Det dataset in Tab. 3. Following the evaluation protocol from GenerateU [18], we compute semantic similarity between generated category embeddings and V3Det category embeddings, selecting the highest similarity match as the final prediction. Our experiments reveal poor performance: GenerateU [18] achieves only 0.4% AP, while the more recent ChatRex [14] achieves just 1.3% AP. This highlights a fundamental limitation: generation-based methods struggle to control the granularity of generated categories, creating semantic misalignments with the specific requirements of detection tasks. Furthermore, even when we finetune GenerateU on V3Det data, the resulting performance (21.8% AP) still exhibits a substantial gap compared to classification-head methods.
Focal Loss Parameters Analysis under Gradient Dilution. As discussed in Sec. 3.1, vast vocabulary detection suffers from gradient dilution challenges. Focal Loss (FL) [30] mitigates this via adaptive weighting with hyperparameters $\alpha$ and $\gamma$. Theoretically, $\alpha$ balances positive/negative sample contributions, and increasing $\alpha$ enhances the model’s ability to handle more categories. The factor $\gamma$ focuses learning on hard negatives, where increasing $\gamma$ improves the mining of hard examples. Tab. 8 evaluates FL configurations on DINO. The default setting [35] ($\alpha=0.25$, $\gamma=2$) achieves 43.4% AP. Tuning these hyperparameters reveals that training is unstable when $\gamma\geq 5$ or $\alpha\geq 0.75$. Notably, the optimal configuration ($\gamma=2$ and $\alpha=0.5$) achieves 47.4% AP, surpassing the baseline by 4.0%. Nonetheless, this is still lower than the 52.3% AP achieved by our CQ-DINO, suggesting that FL hyperparameter tuning, while beneficial, leaves room for further improvement. Interestingly, when examining parameter transferability across architectures in Appendix Tab. 9, we find that this optimal setting does not generalize well. Applying the Swin-B-22k optimal parameters ($\alpha=0.5$, $\gamma=2$) to Swin-B degrades performance by 3.3 % AP relative to its default setting. However, they increase the performance by 4.0% AP and 1.6% AP for Swin-B-22k and Swin-L backbones, respectively. These findings suggest that while Focal Loss is effective in addressing the gradient dilution challenge, optimal hyperparameter selection remains architecture-dependent and requires careful tuning.
4.5 Limitation
Although CQ-DINO improves vast vocabulary object detection, several limitations remain. First, detection performance is influenced by the recall of the category query selection. Fortunately, CQ-DINO achieves 83.3% $AR^{C}$. Appendix Tab. 10 shows that increasing the number of category queries improves $AR^{C}$ but does not lead to higher $AP$. This indicates, in most cases, 83.3% $AR^{C}$ is not a primary bottleneck. Future work will explore more sophisticated selection strategies to address this gap. Second, the two-stage training paradigm, while efficient in practice (first stage requires only $\sim$1 hour), may yield suboptimal coordination between stages compared to end-to-end alternatives.
5 Conclusion
In this work, we systematically analyze the challenges inherent in vast vocabulary detection: positive gradient dilution and hard negative gradient dilution. Through comprehensive experiments, we expose the limitations of Focal Loss under these challenging settings. To mitigate these issues, we propose CQ-DINO, a novel framework with two core innovations: (1) learnable category queries that encode category correlations, and (2) image-guided query selection that effectively reduces the negative space while performing adaptive hard negative mining. Extensive evaluations on the V3Det and COCO benchmarks demonstrate that CQ-DINO achieves superior performance and strong scalability as vocabulary sizes increase. As future work, we plan to investigate the adaptability of our category query formulation for open-vocabulary detection and incremental learning scenarios.
Appendix A Appendix
A.1 Performance of DINO with Different Focal Loss Parameters
In Tab. 8, we achieve the optimal Focal Loss parameters at $\alpha=0.5$ and $\gamma=2$ with the Swin-B-22k backbone. We conduct parameter setting experiments with different backbone in Tab. 9. We find that this optimal setting does not generalize well. The same parameters ($\alpha=0.5$, $\gamma=2$) lead to a performance degradation of 3.3% AP for the Swin-B backbone compared to the default configuration ($\alpha=0.25$, $\gamma=2$). However, they improve the performance by 4.0% AP and 1.6% AP for the Swin-B-22k and Swin-L backbones, respectively. These findings show that while Focal loss addresses gradient dilution issues, its optimal configuration requires careful parameter tuning.
@@ A.1 Performance of DINO with Different Focal Loss Parameters
A.2 Gradient Norm Visualization During Training
As reported in Tab. 8, CQ-DINO achieves superior performance compared to DINO with optimal Focal Loss parameters ($\alpha=0.5$, $\gamma=2$), outperforming the default configuration of DINO ($\alpha=0.25$, $\gamma=2$). To analyze their impact on training gradients, we visualize the gradient norm in Fig. 6 for three configurations: (1) CQ-DINO (red), (2) DINO with $\alpha=0.5$ (blue), and (3) DINO with $\alpha=0.25$ (orange), all using the Swin-B-22k backbone. The results reveal that the gradient norm for DINO with $\alpha=0.25$ (orange line) remains low throughout training, indicating insufficient learning from both positive and hard negative samples. In contrast, DINO with $\alpha=0.5$ (blue line) initially displays strong gradients, but these are unstable and fluctuate considerably, as reflected in the high variance of the blue points. Meanwhile, CQ-DINO maintains a balanced trajectory, sustaining moderate and stable gradient magnitudes during the entire training process. This sustained and balanced gradient norm demonstrates that CQ-DINO more effectively addresses the gradient dilution issues.
Figure 6: Gradient norm visualization during training process.
A.3 Impact of Category Query Count.
Tab. 10 presents the impact of varying the number of category queries (50, 100, 200) on performance. As expected, increasing the number of queries leads to higher $AR^{C}$ scores, at the cost of reduced inference speed. However, simply increasing the query count is not always beneficial. Excessive queries lead to gradient imbalance between positive and negative examples, resulting in diminishing performance gains. Furthermore, improvements in AR do not necessarily correlate with increases in AP, as reflected in our observation that an $AR^{C}$ of 83.3% is not the limiting factor at the current stage. Empirically, we find that 100 queries achieve an optimal balance between detection performance and computational efficiency for V3Det. On the COCO dataset, 30 queries are sufficient, achieving 99.1% $AR^{C}$.
Query Count $AP$ $AR^{C}$ FPS
50 51.8 78.4 6.9
100 52.3 83.3 6.6
200 52.2 87.5 5.9